Vick has 20 meters of fencing to make a rectangular pen for gila monsters what is the largest area that he can use to fence his pet
Answer by amarjeeth123(569)

Perimeter=20 metres
2(length+width)=20
length+width=10
Area is maximum when length and width are maximum and equal.
length=width=10/2=5 metres.
Maximum area=length*width=5(5)=25 square metres.
The maximum area is 25 square metres.